OpenStudy (kalebj):

Derek reads at a rate of 25 words per minute. If he reads constantly until he is finished, how long would it take him to read a 225 word article? A) 93.75 hours B) 10 hours C) 5 hours D) 9 minutes

OpenStudy (kalebj):

I WILL MEDAL AND FAN

OpenStudy (will.h):

25 words per minute means 25x We want ti find how much would it take to read 225 words So we will have this equation 25x = 225 Solve for x

OpenStudy (kalebj):

U have to explain ur answer though ☺ ☺ ☺ ☺ ☺

OpenStudy (kalebj):

thx

OpenStudy (will.h):

You welcome. So ur answer is?

OpenStudy (kalebj):

5 hours

OpenStudy (will.h):

No

OpenStudy (kalebj):

ah... so its 25x = 225 right

OpenStudy (will.h):

25x = 225 To find x (time in minutes) divide both sides by 25 25x/25 = 225/25 X = 9 minutes.

OpenStudy (kalebj):

i got it......9 hours

OpenStudy (will.h):

X represent minutes not hours

OpenStudy (kalebj):

yeah 9 minutes haha

OpenStudy (kalebj):

okay thx im gonna medal and fan now

OpenStudy (will.h):

Logically if he reads 25 words in one minute that means 50 words in 2 minutes. So 225 words would take 5 hrs? Absolutely not

OpenStudy (will.h):

You welcome. I'm just helping people. Medals are not my goal

OpenStudy (kalebj):

well u deserve it thx

OpenStudy (will.h):

My pleasure

OpenStudy (kalebj):

well u deserve it thx

OpenStudy (kalebj):

well u deserve it thx

OpenStudy (kalebj):

well u deserve it thx